Jane,Andre,and Maria pick apples.Andre picks 3 times as many pounds as Maria.Jane picks 2 times as many pounds as Andre.The total weight of apples is 840 pounds.How many pounds of apples does Andre pick?
A 84 pounds
B 252 pounds
C 504 pounds
D 840 pounds

Answer by ptaylor(2198)

Let x=apples that Andre picks
Then x/3=apples that Maria picks
And 2x=apples that Jane picks
Soooo,
x +(x/3)+2x=840 multiply each term by 3
3x+x+6x=2520 collect like terms
10x=2520
x=252#----The amount of apples picked by Andre
x/3 = 252/3=84#---amount of apples picked by Maria
2x=2*252=504#---amount of apples picked by Jane
Hope this helps--ptaylor